# Deep Submillimeter and Radio Observations in the SSA22 Field. III. Multiwavelength Identifications and Properties of the 850 -selected Submillimeter Galaxies

### v1

@@ II.3 Optical/Near-infrared
The optical and near-infrared (opt/NIR) multiband images are centered on the LAEs overdensity peak region (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 107; T. Hayashino, Y. Matsuda, H. Tamura, R. Yamauchi, T. Yamada, M. Ajiki, S. S. Fujita, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2004); Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); T. Yamada, Y. Nakamura, Y. Matsuda, T. Hayashino, R. Yamauchi, N. Morimoto, K. Kousai, and M. Umemura (2012)) (Figure 4). For a detailed description and data processing of the dataset, please refer to Mawatari et al. (2023).
The dataset includes the following multiband data: the CFHT/MegaPrime (Boulade et al., 2003) u*-band data; the Subaru/Suprime-Cam (Miyazaki et al., 2002) B, V, R, i′, z′ band images (Hayashino et al., 2004; Matsuda et al., 2004); the Subaru/HSC (Miyazaki et al., 2012, 2018) g, r, i, z, Y band images collected from the Subaru Strategic Program (HSC-SSP; Aihara et al., 2018); and the Subaru/MOIRCS (Ichikawa et al., 2006) J, H, Ks band data (Uchimoto et al., 2008, 2012). Additionally, the dataset includes the UKIRT/WFCAM (Casali et al., 2007) J, K band data from the deep extragalactic survey, which was part of the UKIRT Infrared Deep Sky Survey (UKIDSS/DXS; Lawrence et al., 2007). The two groups of near-infrared bands are similar, but the footprints of WFCAM are more comprehensive, and the MOIRCS bands are deeper.
@@
In Section III.2, We identified galaxies with extreme red colors: 2 as potential counterparts to SMGs. The color measurement was conducted on PSF-matched images that were smoothed to the largest PSF within the dataset, all having a Full Width at Half Maximum (FWHM) of approximately 1.1. Given the reduced impact of dust extinction in the near-infrared bands, we employed a 2.2 aperture for photometry on the -band images (WFCAM/ and MOIRCS/ bands), while processing the Scam band using the dual mode of SExtractor with identical configuration. To ensure the reliability of this approach, we focused on sources detected at 4 significance in the band, simultaneously reducing the spatial density of potential candidates.
Finally, we estimate the depth of the opt/NIR imaging data. However, on small scales, pixel-to-pixel correlations may arise from either source confusion or systematic biases introduced during data reduction. Similarly, on large scale, pixel correlations can result from imperfect background subtraction, contamination by extended sources, undetected faint sources, or systematic errors such as flat-fielding inaccuracies (Straatman et al., 2016). Hence, we perform aperture photometry 10000 times using a 2 radius on background-subtracted and detected sources masked images in each band, generated by SExtractor. The photometry values represent noise distribution, from which we calculate their standard deviation and convert it to magnitudes as the 1 depth of the images, referred to as the 1 limiting magnitudes (Matsuda et al., 2004; An et al., 2018; Weaver et al., 2022). Based on this, we estimate the corresponding 3 depth of the images for each band, which are summarized in Table 1, along with additional multiband data depths, resolutions.
@@ III COUNTERPART IDENTIFICATION
We primarily utilize four methods, namely VLA 3 GHz, Spitzer/MIPS 24 , Spitzer/IRAC 8 , and optical/NIR red-selected galaxies (hereafter “/radio”,“”, “”, “” metohd), to search for potential counterpart galaxies of SCUBA-2 sources.
The radio and identification methods are the most reliable and widely used approaches for pinpointing SMG counterparts (Barger et al., 2000; Ivison et al., 2002; Umehata et al., 2014; Zavala et al., 2018; Lim et al., 2020b), achieving positional accuracies of 50–80% (Hodge et al., 2013; Chen et al., 2016). However, their completeness is somewhat lower, typically 40–55% 666Notably, Chen et al. (2016) point out that radio identification alone achieves a higher completeness of 80%.. ADF22 (umehata2017; Umehata et al., 2018) contributed 5 additional SMG counterparts (two of which were also identified by the method). In subsequent analyses, we designate sources identified through these methods as ‘MAIN’ samples. To fully exploit the potential of multiwavelength data, we also incorporated and methods (Ashby et al., 2006; Caputi et al., 2006; Schael, 2009; Michałowski et al., 2012; Koprowski, 2015). These approaches offer wider coverage and enable counterparts identification for marginal SCUBA-2 sources, which we classified as ‘SUPP’.
@@
Several studies have selected sole counterparts based on the minimum p-value criterion(Michałowski et al., 2012; Zavala et al., 2018), demonstrating that this approach typically recovers the majority of submillimeter flux(Simpson et al., 2015; Michałowski et al., 2017). However, we adopt a more conservative approach by retaining all potential counterparts.
By combining all identification methods, we matched 248 counterparts to 192 SCUBA-2 sources. The vast majority () of these identified galaxies have deblended 850 flux densities greater than 1 mJy. Notably, for the 212 newly detected SCUBA-2 sources (Section II.1), we successfully identified 136 SMG counterparts for 101 SCUBA-2 sources, representing more than 50% of the total confirmed counterparts. Figure 8 presents representative examples of multiwavelength counterpart identification.
We consider these methods to be generally robust and represent our best-effort approach. However, we emphasize that they are not perfect: some genuine SMGs are inevitably missed, while a fraction of spurious sources may also be included.
@@
| id ID | S aaFor some sources, the deblended fluxes are poorly constrained due to source blending or low signal-to-noise. | id method bbThe 3 count sensitivity limit. | p-value | z-flag ccRedshift types include spectroscopic redshifts, EAZY photometric redshifts, and CIGALE SED-fitting redshifts. When multiple redshift values are available, we adopt the most reliable one according to the hierarchy: (Section IV.2.2). | log | log SFR | log | log | MS ddThe offset of SMGs from the main sequence, defined as MS = log10(sSFR/sSFR), where the main sequence relation is adopted from Speagle et al. (2014). See also Paper IV. | ||
@@
III.1 Radio+ Identifications
@@
Due to the radio-FIR correlation spanning several orders of magnitude in luminosity (Condon, 1992), it remains well-preserved even at high redshifts (Michałowski et al., 2010). Additionally, with its relatively low spatial density (Koprowski, 2015), high-precision radio interferometric observations are widely used for counterpart studies of submillimeter sources (Dudzevičiūtė et al., 2020b). Although the origin of the radio-FIR correlation is debatable, it is widely believed to trace the activity of massive stars () in galaxies (Helou et al., 1985; Condon, 1992; Dudzevičiūtė et al., 2020b). Submillimeter radiation arises from the reemission of ultraviolet photons absorbed by dust from star-forming activities in galaxies (Blain, 2002). Radio emissions in galaxies arise from synchrotron and free-free radiation. Due to their flat spectrum, free-free emissions are weaker than synchrotron emissions at wavelengths longer than 1 cm (frequencies 30 GHz). Therefore, radio emissions are generally attributed to the synchrotron acceleration of relativistic electrons produced by HII regions of massive stars and the supernova remnants left after their death (Condon, 1992).
Similar to the former, the MIR emissions also exhibit a strong correlation with FIR emissions (Appleton et al., 2004; Takeuchi et al., 2005), serving as a reliable indicator of the infrared luminosity of dust-obscured galaxies (Toba et al., 2017; Shim et al., 2022). A significant portion of the MIR emissions from SMGs originates from thermal emissions, generated by dust grains heated by star-forming regions, AGN, or nuclear starburst regions. Moreover, MIR emissions also stem from atomic and molecular lines, primarily induced by the bending and stretching of polycyclic aromatic hydrocarbons (PAHs), excited by UV photons associated with nearby star formation (Menéndez-Delmestre et al., 2009; Casey et al., 2014; Boquien et al., 2019). The dominant PAH feature lines are emitted at rest-frame wavelengths of 3.3, 6.2, 7.7, 8.6, 11.3, 12.7, and 17 µm(Draine and Li, 2007; Dudzevičiūtė et al., 2020b), falling within the observed 24 band at typical redshifts of SMGs.
@@
where n is the surface density of VLA 3 GHz (or 24 ) sources, and is the angular distance between identifications and the SCUBA-2 source positions. For the search radius of submillimeter sources, Hodge et al. (2013) found that, for single-dish maps where blending and confusion are significant, the offset between submillimeter sources and corresponding SMGs does not seem to vary with SNR. Chen et al. (2016) also suggest that using a fixed search radius could be a better strategy, and their test results indicate that adopting 0.1 balances accuracy and completeness. Considering telescope pointing accuracy, errors in source extraction methods, and the impact of source blending, we adopt a fixed search radius of 7 (half of the SCUBA-2 beam size) and designate sources with 0.1 as potential counterparts.
Our VLA 3 GHz and MIPS 24 catalog includes 1989 radio sources with a significance greater than 4 (the image depth after primary beam correction is 1.5-5 Jy), and 813 sources of 3 at 24 , respectively. We cross-matched the SCUBA-2 catalog with them and considered or sources with a significance level of 0.1 as counterparts. For convenience, we will continue the tradition of labeling sources in the catalog as ‘robust’ for 0.05 and ‘extensive’ for 0.05. All identified sources have -values less than 0.1 within 5.5 search radius, while due to the lower density, MIPS identified sources within 6 are considered robust.
@@
After visually inspecting all counterparts and candidate sources, we defined sources with coordinate separations less than 1.5 as the same counterpart. This threshold is appropriate as it avoids misclassifying sources that are too far apart as the same, or splitting the same source into different counterparts. Multiple identifications for the same SCUBA-2 source were sorted in ascending order of p-values and assigned labels a, b, c, and so forth. Additionally, we found that despite the larger FWHM of 24 sources compared to high-precision VLA detections, their coordinates are still relatively accurate.
III.1.1 Identification Result

### v2

@@ II.3 Optical/Near-infrared
The optical and near-infrared (opt/NIR) multi-band images are centered on the Lyman- emitters (LAEs) overdensity peak region (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113; T. Hayashino, Y. Matsuda, H. Tamura, R. Yamauchi, T. Yamada, M. Ajiki, S. S. Fujita, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2004); Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); T. Yamada, Y. Nakamura, Y. Matsuda, T. Hayashino, R. Yamauchi, N. Morimoto, K. Kousai, and M. Umemura (2012)) (Figure 4). For a detailed description and data processing of the dataset, please refer to Mawatari et al. (2023).
The dataset includes the following multi-band data: the CFHT/MegaPrime (Boulade et al., 2003) u*-band data; the Subaru/Suprime-Cam (Miyazaki et al., 2002) B, V, R, i′, z′ band images (Hayashino et al., 2004; Matsuda et al., 2004); the Subaru/HSC (Miyazaki et al., 2012, 2018) g, r, i, z, Y band images collected from the Subaru Strategic Program (HSC-SSP; Aihara et al., 2018); and the Subaru/MOIRCS (Ichikawa et al., 2006) J, H, Ks band data (Uchimoto et al., 2008, 2012). Additionally, the dataset includes the UKIRT/WFCAM (Casali et al., 2007) J, K band data from the deep extragalactic survey, which was part of the UKIRT Infrared Deep Sky Survey (UKIDSS/DXS; Lawrence et al., 2007). The two groups of near-infrared bands are similar, but the footprints of WFCAM are more comprehensive, and the MOIRCS bands are deeper.
@@
In Section III.1.2, We identified galaxies with extreme red colors: 2 as potential counterparts to SMGs. The color measurement was conducted on PSF-matched images that were smoothed to the largest PSF within the dataset, all having a Full Width at Half Maximum (FWHM) of approximately 1.1. Given the reduced impact of dust extinction in the near-infrared bands, we employed a 2.2 aperture for photometry on the -band images (WFCAM/ and MOIRCS/ bands), while processing the Scam band using the dual mode of SExtractor with identical configuration. To ensure the reliability of this approach, we focused on sources detected at 4 significance in the band, simultaneously reducing the spatial density of potential candidates.
Finally, we estimate the depth of the opt/NIR imaging data. However, on small scales, pixel-to-pixel correlations may arise from either source confusion or systematic biases introduced during data reduction. Similarly, on large scale, pixel correlations can result from imperfect background subtraction, contamination by extended sources, undetected faint sources, or systematic errors such as flat-fielding inaccuracies (Straatman et al., 2016). Hence, we perform aperture photometry 10000 times using a 2 radius on background-subtracted and detected sources masked images in each band, generated by SExtractor. The photometry values represent noise distribution, from which we calculate their standard deviation and convert it to magnitudes as the 1 depth of the images, referred to as the 1 limiting magnitudes (Matsuda et al., 2004; An et al., 2018; Weaver et al., 2022). Based on this, we estimate the corresponding 3 depth of the images for each band, which are summarized in Table 1, along with additional multi-band data depths, resolutions.
@@ III COUNTERPART IDENTIFICATION
We primarily utilize four methods, namely VLA 3 GHz, /MIPS 24 , /IRAC 8 , and optical/NIR red-selected galaxies (hereafter “/radio”,“”, “”, “” method), to search for potential counterpart galaxies of SCUBA-2 sources.
The radio and identification methods (Section III.1.1) are the most reliable and widely used approaches for pinpointing SMG counterparts (Barger et al., 2000; Ivison et al., 2002; Umehata et al., 2014; Zavala et al., 2018; Lim et al., 2020b), achieving positional accuracies of 50–80% (Hodge et al., 2013; Chen et al., 2016; An et al., 2018). However, their completeness is somewhat lower, typically 40–55% 666Notably, Chen et al. (2016) point out that radio identification alone achieves a higher completeness of 80%.. ADF22 (Umehata et al., 2017, 2018) contributed 5 additional SMG counterparts (two of which were also identified by the method). In subsequent analyses, we designate sources identified through these methods as ‘MAIN’ samples. To fully exploit the potential of multiwavelength data, we also incorporated and methods (Ashby et al., 2006; Caputi et al., 2006; Schael, 2009; Koprowski, 2015) (Section III.1.2). These approaches offer wider coverage and enable counterparts identification for marginal SCUBA-2 sources, which we classified as ‘SUPP’. Section III.1 outlines the feasibility and reliability of each identification method.
@@
Several studies have selected sole counterparts based on the minimum -value criterion (Zavala et al., 2018), demonstrating that this approach typically recovers the majority of submillimeter flux (Simpson et al., 2015). However, we adopt a more conservative approach by retaining all potential counterparts.
By combining all identification methods, we matched 248 counterparts to 192 SCUBA-2 sources. The vast majority () of these identified galaxies have deblended 850 flux densities greater than 1 mJy. Notably, for the 212 newly detected SCUBA-2 sources (Section II.1), we successfully identified 136 SMG counterparts for 101 SCUBA-2 sources, representing more than 50% of the total confirmed counterparts. Figure 8 presents representative examples of multiwavelength counterpart identification. A more detailed breakdown of the identifications is provided in Section III.2.
We consider these methods to be generally robust and represent our best-effort approach. However, we emphasize that they are not perfect: some genuine SMGs are inevitably missed, while a fraction of spurious sources may also be included. We offer an in-depth discussion of the reliability of these identification methods and their potential limitations in Section III.3.
@@
| id ID | S aaFor some sources, the deblended fluxes are poorly constrained due to source blending or low signal-to-noise. | id method bbThe 3 count sensitivity limit. | -value | z-flag ccRedshift types include spectroscopic redshifts, EAZY photometric redshifts, and CIGALE SED-fitting redshifts. When multiple redshift values are available, we adopt the most reliable one according to the hierarchy: (Section IV.2.2). | log | log SFR | log | log | MS ddThe offset of SMGs from the main sequence, defined as MS = log10(sSFR/sSFR), where the main sequence relation is adopted from Speagle et al. (2014). See also Paper IV. | ||
@@
III.1 Identification Methods
III.1.1 Radio+ Methods
@@
Due to the radio-FIR correlation spanning several orders of magnitude in luminosity (Condon, 1992), it remains well-preserved even at high redshifts. Additionally, with its relatively low spatial density (Koprowski, 2015), high-precision radio interferometric observations are widely used for counterpart studies of submillimeter sources (Dudzevičiūtė et al., 2020b). Although the origin of the radio-FIR correlation is debatable, it is widely believed to trace the activity of massive stars () in galaxies (Helou et al., 1985; Condon, 1992; Dudzevičiūtė et al., 2020b). Submillimeter radiation arises from the re-emission of ultraviolet photons absorbed by dust from star-forming activities in galaxies (Blain, 2002). Radio emissions in galaxies arise from synchrotron and free-free radiation. Due to their flat spectrum, free-free emissions are weaker than synchrotron emissions at wavelengths longer than 1 cm (frequencies 30 GHz). Therefore, radio emissions are generally attributed to the synchrotron acceleration of relativistic electrons produced by HII regions of massive stars and the supernova remnants left after their death (Condon, 1992).
Similar to the former, the MIR emissions also exhibit a strong correlation with FIR emissions (Appleton et al., 2004; Takeuchi et al., 2005), serving as a reliable indicator of the infrared luminosity of dust-obscured galaxies (Toba et al., 2017; Shim et al., 2022). A significant portion of the MIR emissions from SMGs originates from thermal emissions, generated by dust grains heated by star-forming regions, AGN, or nuclear starburst regions. Moreover, MIR emissions also stem from atomic and molecular lines, primarily induced by the bending and stretching of polycyclic aromatic hydrocarbons (PAHs), excited by UV photons associated with nearby star formation (Menéndez-Delmestre et al., 2009; Casey et al., 2014; Boquien et al., 2019). The dominant PAH feature lines are emitted at rest-frame wavelengths of 3.3, 6.2, 7.7, 8.6, 11.3, 12.7, and 17 (Draine and Li, 2007; Dudzevičiūtė et al., 2020b), falling within the observed 24 band at typical redshifts of SMGs. We should note that it is precisely due to these prominent PAH emission and absorption features (Huang et al., 2009) that the 24 flux in SMGs can exhibit significant variations.
@@
where is the surface density of VLA 3 GHz (or 24 ) sources, and is the angular distance between identifications and the SCUBA-2 source positions.
Hodge et al. (2013) argued that, even in single-dish submillimeter maps where source confusion is significant, the positional offset between the submillimeter source and true SMG counterpart does not appear to vary systematically with SNR. An et al. (2018) adopted a statistical threshold of 0.065 for reliable radio identifications, though they noted that a small fraction of confirmed radio counterparts still lie beyond this threshold—corresponding to 5.5. Similarly, Chen et al. (2016) found that using a fixed search radius for counterpart identification could be a better strategy, and suggested that adopting 0.1 balances accuracy and completeness.
We therefore adopt a fixed search radius of 7 (half of the SCUBA-2 beam size) and consider all VLA and MIPS 24 counterparts with 0.1 as potential associations. Specifically, all VLA sources within a 5.4 radius of a SCUBA-2 centroid have 0.1. For example, in Figure 9, the sources SSA22.0000, SSA22.0070, and SSA22.0002 are affected by source blending. Our identification method did not associate the radio counterparts located near the edge of the SCUBA-2 beam with these sources; however, ALMA detections confirm that these radio sources are indeed physically associated SMGs. For MIPS 24 , due to its lower source surface density, all counterparts within the full SCUBA-2 beam (7) satisfy 0.065. Taking into account SCUBA-2 beam effect, source blending, uncertainties in source extraction, astrometric errors, and telescope pointing accuracy ( 1–2 for JCMT; Geach et al., 2017), we regard the 0.1 criterion as well justified for radio counterparts. Given that 24 band has a resolution of 6, we similarly consider this threshold appropriate for 24 identifications. For the IRAC 8 and -color identifications used in the next section, we adopt a stricter threshold of 0.05, corresponding to angular radii of approximately 2.4 and 2.8, respectively. We provide a more detailed justification for these choices in Section III.3.
@@
For convenience, we will continue the tradition of labeling sources in the catalog as ‘robust’ for 0.05 and ‘tentative’ for 0.05. After visually inspecting all counterparts and candidate sources, we defined sources with coordinate separations less than 1.5 as the same counterpart. This threshold is appropriate as it avoids misclassifying sources that are too far apart as the same, or splitting the same source into different counterparts. Multiple identifications for the same SCUBA-2 source were sorted in ascending order of -values and assigned labels a, b, c, and so forth. Additionally, we found that despite the larger FWHM of 24 sources compared to high-precision VLA detections, their coordinates are still relatively accurate.
III.1.2 + Methods
We also employ IRAC 8 and red opt-NIR color as a method to search for potential counterparts. However, for reliability considerations, we only include those counterparts with 0.05 in the source catalog.
Similar to the MIR-FIR correlation mentioned earlier, exploration in the IRAC 8 band is often conducted to enhance the completeness of identifying submillimeter sources (Ashby et al., 2006; Koprowski, 2015; Koprowski et al., 2017; Zavala et al., 2018). SED analysis of SMGs suggested that their 8 emission primarily traces ongoing star formation (Ashby et al., 2006). The evolved stellar populations have been discovered in massive SMGs, the expected spectral signatures of evolved stellar populations should peak at rest-frame 2-3 (Koprowski, 2015), corresponding to observed 8 .
Some studies have used a NIR color to select high-redshift, extremely red galaxies, such as HIEROs (Wang et al., 2012), KIEROs (Wang et al., 2016), OIRTC (Chen et al., 2016). Similarly, due to the highly dust-obscured nature of SMGs often displaying red optical/NIR colors, many researchers have used color-selected galaxies to identify DSFGs or SMGs (Cowie et al., 2017; An et al., 2018; Zhang et al., 2022; McKay et al., 2025). Following Michałowski et al. (2012), we employed a red optical/NIR color criteria ( 2 & 23 mag) to identify potential SMGs counterparts. The detailed measurement description could be seen in the Section II.3.
SMGs typically exhibit redshifts in the range –3. In this redshift interval, the observed - and -bands correspond approximately to the rest-frame near-UV and -band, respectively, straddling the Balmer/4000Å break—a key spectral feature for galaxy characterization. The Balmer break arises at the Balmer limit (3645 Å) due to bound-free absorption by hydrogen atoms with electrons in the energy level. This feature is weakened in hot OB stars because most hydrogen is ionized, but it is strongest in cooler A-type stars (– M⊙), whose effective temperatures (10,000 K) maximize the population of electrons in the state. This feature is particularly prominent in star-forming galaxies (Dudzevičiūtė et al., 2020b; Wilkins et al., 2023). Moreover, it becomes especially strong in galaxies with stellar populations of age 0.3 Gyr, when a large number of UV-bright stars begin to evolve off the main sequence (Bruzual and Charlot, 2003; Vikaeus et al., 2024). The 4000 Å break primarily driven by metal absorption lines from cool, evolved stars (Mo et al., 2010; Dudzevičiūtė et al., 2020b). Together, the Balmer/4000Å break serves as a powerful diagnostic of recent star formation history and is sensitive to dust attenuation, metallicity, and the Lyman continuum (LyC) escape fraction (Wilkins et al., 2023).
These selected galaxies are referred to as Extremely Red Galaxies (ERGs), which may stem from evolved stellar components or galaxies that have reddened significantly due to dust obscuration in star-forming galaxies or starbursts (or AGN) (Caputi et al., 2006; Schael, 2009; Mo et al., 2010). ERGs are believed to align on an evolutionary sequence with high-redshift SMGs and local massive ellipticals (Smail et al., 2002; Schael, 2009), while the most extreme ERGs represent a population of highly dust-obscured, ultraluminous galaxies at high redshift (Smail et al., 1999). In the redshift range of 1.5-3, 40-60% of high-mass ERGs are ULIRGs (Smail et al., 2002; Caputi et al., 2006). SMGs, characterized by severe dust obscuration and intense star formation, are expected to exhibit distinct optical/NIR features. Therefore, such selections seem to be a good criterion.
III.2 Identification Result
Our VLA 3 GHz and MIPS 24 catalog includes 1989 radio sources with a significance greater than 4 (the image depth after primary beam correction is 1.5-5 Jy), and 813 sources of 3 at 24 , respectively.